# DG Singular equivalence and singular locus

### v2

@@ 1.1. Background
Let $A$ be an associative algebra over a base field $k$ of characteristic zero. Its singularity category $D_{sg}(A)$ is the Verdier quotient $D^{b}(A)/\mathrm{Perf}(A)$, where $\mathrm{Perf}(A)$ is the full subcategory consisting of perfect complexes over $A$. $D_{sg}(A)$ measures the smoothness of $A$ in the sense that $A$ is homologically smooth if and only if $D_{sg}(A)$ is trivial. It was first introduced by Buchweitz in the study of algebraic representations of Gorenstein rings in [5]. Moreover, he showed that $D_{sg}(A)$ is equivalent to the Cohen-Macaulay stable category $\underline{\mathrm{CM}}(A)$ as triangulated categories when $A$ is Gorenstein. Later, in [20], Orlov rediscovered this notion from the perspective of algebraic geometry and mathematical physics, which has a deep relationship with Homological Mirror Symmetry.
@@ 1.2. Main result
Let $R$ be a commutative Gorenstein Noetherian ring. Let $S_{dg}(R)$ be the triangulated category described as above. Then there is an affine scheme $X$ constructed solely from $S_{dg}(R)$, and a finite surjective morphism $X\rightarrow\mathrm{Spec}(R/I)$, where $\mathrm{Spec}(R/I)$ is the singular locus in $\mathrm{Spec}(R)$.
@@ 2.1. Singularity category and singular equivalence
    for any $M^{\ast},N^{\ast}\in\mathrm{Ob}\big{(}D_{sg}(A)\big{)}$ and $f\in\mathrm{Hom}_{D^{b}(A)}(M^{\ast},N^{\ast})$ such that $\mathrm{Cone}(f)\in\mathrm{Ob}\big{(}\mathrm{Perf}(A)\big{)}$ and $f$ is invertible in $D_{sg}(A)$. Denote by $\mathcal{S}$ the set of morphisms satisfying the above condition in $D^{b}(A)$. Then the homomorphism space $\mathrm{Hom}_{D_{sg}(A)}(M^{*},N^{*})$ of $D_{sg}(A)$ is the localization of $\mathrm{Hom}_{D^{b}(A)}(M^{*},N^{*})$ over $\mathcal{S}$.
@@
Let $S:=k[x_{1},\cdots,x_{m}]$ for some integer $m$ and $1\neq f\in S$ be a nontrivial element. Consider two algebras $S[u]/(f)$ and $S[u,v]/(f+uv)$ by adding variables $u$ and $u,v$ respectively. There are two algebra morphisms
$S\hookrightarrow S[u]/(f)$
@@ 3.1. Generalized Tate-Hochschild complex of associative algebra
whose cohomologies, in the case of $\Lambda$ being self-injective, are exactly the Tate-Hochschild cohomologies, where $\Lambda^{\vee}:=\mathrm{Hom}_{\Lambda^{e}}(\Lambda,\Lambda\otimes\Lambda)$, the differential $\mu$ is given by the multiplication of $\Lambda$: $\mu(x\otimes y)=xy$, $C_{\ast}(\Lambda,\Lambda^{\vee})$ is the Hochschild chain complex of $\Lambda$ for $\Lambda^{e}$-module $\Lambda^{\vee}$, and $C^{\ast}(\Lambda,\Lambda)$ is the Hochschild cochain complex of $\Lambda$.
@@
$(-)\otimes^{\mathbb{L}}_{A}M:D_{sg}(S)\rightarrow D_{sg}\big{(}S[u,v]/(f+uv)\big{)}.$
@@
$S_{dg}(S)\simeq S_{dg}\big{(}S[u,v]/(f+uv)\big{)}.$
In [16, Theorem 1.1], Keller realized the Tate-Hochschild cohomologies of an algebra $A$ as the Hochschild cohomologies of $S_{dg}(A)$ in the following theorem.
@@
sending the unique object of $A^{op}$ to $A$-module complex $A$. By the result of Keller in [15, Theorem 8.1], the triangle functor
@@
By the definition of tensor product of DG modules over a DG category (see [18, 3.5]), we know that
@@
If we replace $N^{\ast}$ by its projective $A$-module resolution, say, $P_{N^{*}}$, then by [18, 3.5], $h_{N^{\ast}}\otimes_{\mathcal{P}\mathrm{erf}(A)}^{\mathbb{L}}\tilde{h}_{M^{\ast}}$ is isomorphic to $h_{P_{N^{\ast}}}\otimes_{\mathcal{P}\mathrm{erf}(A)}\tilde{h}_{M^{\ast}}$. It follows that
@@
given by $\mathrm{Bar}(R)\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$. View this double complex as a cochain complex, and denote it by $E^{\ast}$. Also, denote by $F^{\ast}$ the cochain complex obtained from the double complex $\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),\mathrm{Bar}(R))$. We get a complex $\mathrm{Cone}(E^{\ast}\rightarrow F^{\ast})$ and obtain the following proposition for this complex.
@@ 3.2. Hochschild cohomology of DG singularity category
$E^{\ast}\otimes E^{\ast}\cong$ $\Big{(}\mathrm{Bar}(R)\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}\Big{)}\otimes\Big{(}\mathrm{Bar}(R)\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}\Big{)}$
$\cong\mathrm{Bar}(R)\otimes_{R^{e}}\Big{(}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}\otimes\mathrm{Bar}(R)\Big{)}\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$
$\xrightarrow{invol.}\mathrm{Bar}(R)\otimes_{R^{e}}R^{e}\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$
$\cong\mathrm{Bar}(R)\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}=E^{\ast},$
induced by involution, gives a natural multiplication structure on $E^{\ast}$. Finally, there is a canonical $\Big{(}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),\mathrm{Bar}(R)\big{)}\Big{)}^{e}$-module structure on $\mathrm{Bar}(R)\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$. Therefore, viewing $\mathrm{Cone}(E^{\ast}\rightarrow F^{*})$ as a semi-product $F^{\ast}\ltimes E^{\ast}$, we obtain on $\mathrm{Cone}(E^{\ast}\rightarrow F^{*})$ a natural multiplication structure.
@@
$\mathrm{Bar}(R)\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$ $\cong\mathrm{Hom}_{R^{e}}\big{(}R^{e},\mathrm{Bar}(R)\big{)}\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$
@@
We next recall the support schemes of object in $S_{sg}(R)$ and of $S_{sg}(R)$.
@@
(1) Let $Q^{\ast}$ be an object in $S_{sg}(R)$. The support scheme $\mathrm{Supp}(Q^{\ast})$ of $Q^{\ast}$ is a subscheme of $\mathrm{Spec}(R)$:
$\mathrm{Supp}(Q^{\ast}):=\big{\{}\mathbf{p}\in\mathrm{Spec}(R)\mid Q^{\ast}_{\mathbf{p}}\ncong 0\,\textup{in}\,D_{sg}(R_{\mathbf{p}})\big{\}}.$
(2) The support scheme $\mathrm{Supp}(S_{dg}(R))$ of $S_{dg}(R)$ is a subscheme of $\mathrm{Spec}(R)$ :
$\mathrm{Supp}(S_{dg}(R)):=\bigcap\limits_{Q^{\ast}\in\mathrm{Obj}\big{(}S_{dg}(R)\big{)}}\mathrm{Supp}(Q^{\ast}).$
@@
Since $\bigoplus\limits_{i}\mathrm{Hom}_{D_{sg}(R^{e})}(R,R[i])$ contains the unit $\mathrm{id}$ and $\pi_{R}(1_{R})=\mathrm{id}$, its annihilator, as an $R$-module, is trivial. Hence, when we view $\bigoplus\limits_{i}\mathrm{Hom}_{D_{sg}(R^{e})}(R,R[i])$ as a sheaf on $\mathrm{Spec}(R)$, to prove the proposition, it suffices to show that the support scheme of $\bigoplus\limits_{i}\mathrm{Hom}_{D_{sg}(R^{e})}(R,R[i])$ is $V(I)\subseteq\mathrm{Spec}(R)$, which is equivalent to showing that the support scheme of $\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)$ in $D(k)$ is exactly $V(I)$.
@@
$E^{*}_{\mathbf{p}}$ $=\left(\mathrm{Bar}(R)\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}\right)_{\mathbf{p}}$
$\cong\left(R\otimes_{R^{e}}\overline{\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}}\right)_{\mathbf{p}}$
$\cong R_{\mathbf{p}}\otimes_{R^{e}}\overline{\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}}$
$\cong R_{\mathbf{p}}\otimes_{R_{\mathbf{p}}^{e}}R_{\mathbf{p}}^{e}\otimes_{R^{e}}\overline{\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}}$
$\cong(R_{\mathbf{p}}\otimes_{R_{\mathbf{p}}^{e}}R_{\mathbf{p}}^{e})\otimes^{\mathbb{L}}_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$
$\cong(R_{\mathbf{p}}\otimes^{\mathbb{L}}_{R_{\mathbf{p}}^{e}}R_{\mathbf{p}}^{e})\otimes^{\mathbb{L}}_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$
$\cong\mathrm{Bar}(R_{\mathbf{p}})\otimes^{\mathbb{L}}_{R_{\mathbf{p}}^{e}}R_{\mathbf{p}}^{e}\otimes^{\mathbb{L}}_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$
$\cong\mathrm{Bar}(R_{\mathbf{p}})\otimes^{\mathbb{L}}_{R_{\mathbf{p}}^{e}}\Big{(}R_{\mathbf{p}}^{e}\otimes_{R^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}\Big{)}$
$\cong\mathrm{Bar}(R_{\mathbf{p}})\otimes^{\mathbb{L}}_{R_{\mathbf{p}}^{e}}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R_{\mathbf{p}}^{e}\big{)}$
$\cong\mathrm{Bar}(R_{\mathbf{p}})\otimes^{\mathbb{L}}_{R_{\mathbf{p}}^{e}}\mathrm{Hom}_{R_{\mathbf{p}}^{e}}\big{(}\mathrm{Bar}(R)\otimes_{R^{e}}R_{\mathbf{p}}^{e},R_{\mathbf{p}}^{e}\big{)}$
$\cong\mathrm{Bar}(R_{\mathbf{p}})\otimes^{\mathbb{L}}_{R_{\mathbf{p}}^{e}}\mathrm{Hom}_{R_{\mathbf{p}}^{e}}\big{(}\mathrm{Bar}(R_{\mathbf{p}}),R_{\mathbf{p}}^{e}\big{)}$
in $D(k)$, where $\overline{\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}}$ is a flat resolution of $\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$ over $R^{e}$. In the meantime, when we replace $R$ by $R_{\mathbf{q}}$, we obtain that
@@
Now we assume that $\mathbf{p}$ is a non-singular point. Since ${\mathbf{p}}$ is non-singular, $R_{\mathbf{p}}$ is a homologically smooth algebra. Then $\mathrm{Hom}_{S_{dg}(R_{\mathbf{p}}^{e})}(R_{\mathbf{p}},R_{\mathbf{p}})$ is a trivial in $D(k)$. Hence $\left(\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)\right)_{\mathbf{p}}$ is also a trivial in $D(k)$. Thus, ${\mathbf{p}}$ is not contained in the support scheme $\mathrm{Supp}(S_{dg}(R))$ of $\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)$ in $D(k)$, which implies that
$\mathrm{Supp}(S_{dg}(R))\subseteq V(I):=\mathrm{Spec}(R/I).$
@@
Assume that $\mathrm{Hom}_{S_{dg}(R_{\mathbf{q}}^{e})}(R_{\mathbf{q}},R_{\mathbf{q}})$ is a trivial in $D(k)$. It implies that $\mathrm{HH}^{0}_{sg}(R_{\mathbf{q}})$ is a trivial algebra. From the distinguished triangle
@@
is a trivial in $D(k)$ does not hold. Thus $\mathrm{Hom}_{S_{dg}(R_{\mathbf{q}}^{e})}(R_{\mathbf{q}},R_{\mathbf{q}})$ is a nontrivial in $D(k)$. From the isomorphism
$\mathrm{Hom}_{D^{b}(R_{\mathbf{q}}^{e})}(R_{\mathbf{q}},R_{\mathbf{q}})\cong\left(\mathrm{Hom}_{D^{b}(R^{e})}(R,R)\right)_{\mathbf{q}},$
as algebras in the above argument, we know that $\mathbf{q}$ is contained in $\mathrm{Supp}(S_{dg}(R))$, which suggests that
$V(I)\subseteq\mathrm{Supp}(S_{dg}(R)).$
Thus we obtain $V(I)=\mathrm{Supp}(S_{dg}(R))$, that is, the support scheme of $\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)$ in $D(k)$ is $V(I)$. We thus completed the proof. ∎
@@
Second, since $R$ is a Gorenstein Noetherian algebra, $R^{e}$ is also a Gorenstein Noetherian algebra. Hence, $R^{e}$ admits a bounded finitely generated injective $R^{e}$-module resolution, denoted by $J^{*}$. There is a quasi-isomorphism of $R^{e}$-modules complexes:
$\mathbb{R}\mathrm{Hom}_{R}(\mathrm{Bar}(R),J^{*})\rightarrow\mathbb{R}\mathrm{Hom}_{R}(\mathrm{Bar}(R),R^{e})$

### v5

@@ 1.1. Background
Let $A$ be an associative algebra over a base field $k$ of characteristic zero. Its singularity category $D_{sg}(A)$ is the Verdier quotient $D^{b}(A)/\mathrm{Perf}(A)$, where $\mathrm{Perf}(A)$ is the full subcategory consisting of perfect complexes over $A$. $D_{sg}(A)$ measures the smoothness of $A$ in the sense that $A$ is homologically smooth if and only if $D_{sg}(A)$ is trivial. It was first introduced by Buchweitz in the study of algebraic representations of Gorenstein rings in [5]. Moreover, he showed that $D_{sg}(A)$ is equivalent to the stable category $\underline{\mathrm{CM}}(A)$ of Cohen-Macaulay $A$-modules as triangulated categories when $A$ is Gorenstein. Later, in [21], Orlov rediscovered this notion from the perspective of algebraic geometry and mathematical physics, which has a deep relationship with Homological Mirror Symmetry.
@@ 1.2. Main result
Let $R$ be a commutative Gorenstein Noetherian ring. Let $S_{dg}(R)$ be the DG category described as above. Then there is an affine scheme $X$ constructed solely from $S_{dg}(R)$, and a finite surjective morphism $X\rightarrow\mathrm{Spec}(R/I)$, where $\mathrm{Spec}(R/I)$ is the singular locus in $\mathrm{Spec}(R)$.
@@ 2.1. Singularity category and singular equivalence
    for any $M^{\ast},N^{\ast}\in\mathrm{Ob}\big{(}D_{sg}(A)\big{)}$ and $f\in\mathrm{Hom}_{D^{b}(A)}(M^{\ast},N^{\ast})$ such that $\mathrm{Cone}(f)\in\mathrm{Ob}\big{(}\mathrm{Perf}(A)\big{)}$, let $f$ be invertible in $D_{sg}(A)$. Denote by $\mathcal{S}$ the set of morphisms satisfying the above condition in $D^{b}(A)$. Then the homomorphism space $\mathrm{Hom}_{D_{sg}(A)}(M^{*},N^{*})$ of $D_{sg}(A)$ is the localization of $\mathrm{Hom}_{D^{b}(A)}(M^{*},N^{*})$ over $\mathcal{S}$.
@@
Let $S:=k[x_{1},\cdots,x_{m}]$ for some integer $m$ and $0\neq f\in S$ be nontrivial element. Consider two algebras $S[u]/(f)$ and $S[u,v]/(f+uv)$ by adding variables $u$ and $u,v$ respectively. There are two algebra morphisms
$S/(f)\hookrightarrow S[u]/(f)$
@@ 3.1. Generalized Tate-Hochschild complex of associative algebra
whose cohomologies, in the case of $\Lambda$ being self-injective, are exactly the Tate-Hochschild cohomologies, where $\Lambda^{\vee}:=\mathrm{Hom}_{\Lambda^{e}}(\Lambda,\Lambda\otimes\Lambda)$, the differential $\mu$ is given by the multiplication of $\Lambda$: $\mu(x\otimes y)=xy$, $C_{\ast}(\Lambda,\Lambda^{\vee})$ is the Hochschild chain complex of $\Lambda$ with $\Lambda^{e}$-module $\Lambda^{\vee}$, and $C^{\ast}(\Lambda,\Lambda)$ is the Hochschild cochain complex of $\Lambda$.
@@
$(-)\otimes^{\mathbb{L}}_{S/(f)}M:D_{sg}(S/(f))\rightarrow D_{sg}\big{(}S[u,v]/(f+uv)\big{)}.$
@@
$S_{dg}(S/(f))\simeq S_{dg}\big{(}S[u,v]/(f+uv)\big{)}.$
In [17, Theorem 1.1], Keller realized the Tate-Hochschild cohomologies of an algebra $A$ as the Hochschild cohomologies of $S_{dg}(A)$ in the following theorem.
@@
sending the unique object of $A^{op}$ to $A$-module complex $A$. By the result of Keller in [16, Theorem 8.1], the triangle functor
@@
By the definition of tensor product of DG modules over a DG category (see [19, 3.5]), we know that
@@
If we replace $N^{\ast}$ by its projective $A$-module resolution, say, $P_{N^{*}}$, then by [19, 3.5], $h_{N^{\ast}}\otimes_{\mathcal{P}\mathrm{erf}(A)}^{\mathbb{L}}\tilde{h}_{M^{\ast}}$ is isomorphic to $h_{P_{N^{\ast}}}\otimes_{\mathcal{P}\mathrm{erf}(A)}\tilde{h}_{M^{\ast}}$. It follows that
@@
given by $\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$. View this double complex as a cochain complex, and denote it by $E^{\ast}$. Also, denote by $F^{\ast}$ the cochain complex obtained from the double complex $\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),\mathrm{Bar}(R))$. We get a complex $\mathrm{Cone}(E^{\ast}\rightarrow F^{\ast})$ and obtain the following proposition for this complex.
@@ 3.2. Hochschild cohomology of DG singularity category
$E^{\ast}\otimes E^{\ast}\cong$ $\Big{(}\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}\Big{)}\otimes\Big{(}\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}\Big{)}$
$\cong\mathrm{Bar}(R)\otimes_{R^{e}}\Big{(}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}\otimes\mathrm{Bar}(R)\Big{)}\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$
$\xrightarrow{invol.}\mathrm{Bar}(R)\otimes_{R^{e}}R^{e}\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$
$\cong\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}=E^{\ast},$
induced by involution, gives a natural multiplication structure on $E^{\ast}$. Finally, there is a canonical $\Big{(}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),\mathrm{Bar}(R)\big{)}\Big{)}^{e}$-module structure on $\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$. Therefore, viewing $\mathrm{Cone}(E^{\ast}\rightarrow F^{*})$ as a semi-product $F^{\ast}\ltimes E^{\ast}$, we obtain on $\mathrm{Cone}(E^{\ast}\rightarrow F^{*})$ a natural multiplication structure.
@@
$\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$ $\cong\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}R^{e},\mathrm{Bar}(R)\big{)}\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$
@@
We next recall the support scheme of object in $D(R)$ and introduce the notion of the diagonal support scheme of $S_{sg}(R)$.
@@
(1) Let $Q^{\ast}$ be an object in $D(R)$. The support scheme $\mathrm{Supp}(Q^{\ast})$ of $Q^{\ast}$ is a subscheme of $\mathrm{Spec}(R)$:
$\mathrm{Supp}(Q^{\ast}):=\big{\{}\mathbf{p}\in\mathrm{Spec}(R)\mid Q^{\ast}_{\mathbf{p}}\ncong 0\,\textup{in}\,D(R_{\mathbf{p}})\big{\}}.$
(2) The diagonal support scheme $\mathrm{DSupp}(S_{dg}(R))$ of $S_{dg}(R)$ is a subscheme of $\mathrm{Spec}(R)$:
$\mathrm{DSupp}(S_{dg}(R)):=\big{\{}\mathbf{p}\in\mathrm{Spec}(R)\mid\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)_{\mathbf{p}}\ncong 0\,\textup{in}\,D(R_{\mathbf{p}})\big{\}}.$
@@
Since $\bigoplus\limits_{i}\mathrm{Hom}_{D_{sg}(R^{e})}(R,R[i])$ contains the unit $\mathrm{id}$ and $\pi_{R}(1_{R})=\mathrm{id}$, its annihilator, as an $R$-module, is trivial. Hence, when we view $\bigoplus\limits_{i}\mathrm{Hom}_{D_{sg}(R^{e})}(R,R[i])$ as a sheaf on $\mathrm{Spec}(R)$, to prove the proposition, it suffices to show that the support scheme of $\bigoplus\limits_{i}\mathrm{Hom}_{D_{sg}(R^{e})}(R,R[i])$ is $V(I)\subseteq\mathrm{Spec}(R)$, which is equivalent to showing that the diagonal support scheme $\mathrm{DSupp}(S_{dg}(R))$ is exactly $V(I)$.
@@
$E^{*}_{\mathbf{p}}$ $=\left(\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}\right)_{\mathbf{p}}$
$\cong\left(R\otimes_{R^{e}}\overline{\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}}\right)_{\mathbf{p}}$
$\cong R_{\mathbf{p}}\otimes_{R^{e}}\overline{\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}}$
$\cong R_{\mathbf{p}}\otimes_{R_{\mathbf{p}}^{e}}R_{\mathbf{p}}^{e}\otimes_{R^{e}}\overline{\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}}$
$\cong(R_{\mathbf{p}}\otimes_{R_{\mathbf{p}}^{e}}R_{\mathbf{p}}^{e})\otimes^{\mathbb{L}}_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$
$\cong(R_{\mathbf{p}}\otimes^{\mathbb{L}}_{R_{\mathbf{p}}^{e}}R_{\mathbf{p}}^{e})\otimes^{\mathbb{L}}_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$
$\cong\mathrm{Bar}(R_{\mathbf{p}})\otimes^{\mathbb{L}}_{R_{\mathbf{p}}^{e}}R_{\mathbf{p}}^{e}\otimes^{\mathbb{L}}_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$
$\cong\mathrm{Bar}(R_{\mathbf{p}})\otimes^{\mathbb{L}}_{R_{\mathbf{p}}^{e}}\Big{(}R_{\mathbf{p}}^{e}\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}\Big{)}$
$\cong\mathrm{Bar}(R_{\mathbf{p}})\otimes^{\mathbb{L}}_{R_{\mathbf{p}}^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R_{\mathbf{p}}^{e}\big{)}$
$\cong\mathrm{Bar}(R_{\mathbf{p}})\otimes^{\mathbb{L}}_{R_{\mathbf{p}}^{e}}\mathbb{R}\mathrm{Hom}_{R_{\mathbf{p}}^{e}}\big{(}\mathrm{Bar}(R)\otimes_{R^{e}}R_{\mathbf{p}}^{e},R_{\mathbf{p}}^{e}\big{)}$
$\cong\mathrm{Bar}(R_{\mathbf{p}})\otimes^{\mathbb{L}}_{R_{\mathbf{p}}^{e}}\mathbb{R}\mathrm{Hom}_{R_{\mathbf{p}}^{e}}\big{(}\mathrm{Bar}(R_{\mathbf{p}}),R_{\mathbf{p}}^{e}\big{)}$
in $D(k)$, where $\overline{\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}}$ is a flat resolution of $\mathbb{R}\mathrm{Hom}_{R^{e}}\big{(}\mathrm{Bar}(R),R^{e}\big{)}$ over $R^{e}$. In the meantime, when we replace $R$ by $R_{\mathbf{p}}$, we obtain that
@@
Now we assume that $\mathbf{p}$ is a non-singular point. Since ${\mathbf{p}}$ is non-singular, $R_{\mathbf{p}}$ is a homologically smooth algebra. Then $\mathrm{Hom}_{S_{dg}(R_{\mathbf{p}}^{e})}(R_{\mathbf{p}},R_{\mathbf{p}})$ is trivial in $D(R_{\mathbf{p}})$. Hence $\left(\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)\right)_{\mathbf{p}}$ is also trivial in $D(R_{\mathbf{p}})$. Thus, ${\mathbf{p}}$ is not contained in the diagonal support scheme $\mathrm{DSupp}(S_{dg}(R))$, which implies that
$\mathrm{DSupp}(S_{dg}(R))\subseteq V(I):=\mathrm{Spec}(R/I).$
@@
Assume that $\mathrm{Hom}_{S_{dg}(R_{\mathbf{q}}^{e})}(R_{\mathbf{q}},R_{\mathbf{q}})$ is trivial in $D(R_{\mathbf{q}})$. It implies that $\mathrm{HH}^{0}_{sg}(R_{\mathbf{q}})$ is trivial algebra. From the distinguished triangle
@@
is trivial in $D(R_{\mathbf{q}})$ does not hold. Thus $\mathrm{Hom}_{S_{dg}(R_{\mathbf{q}}^{e})}(R_{\mathbf{q}},R_{\mathbf{q}})$ is nontrivial in $D(R_{\mathbf{q}})$. From the isomorphism
$\left(\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)\right)_{\mathbf{q}}\cong\mathrm{Hom}_{S_{dg}(R_{\mathbf{q}}^{e})}(R_{\mathbf{q}},R_{\mathbf{q}}),$
as algebras in the above argument, we know that $\mathbf{q}$ is contained in $\mathrm{DSupp}(S_{dg}(R))$, which suggests that
$V(I)\subseteq\mathrm{DSupp}(S_{dg}(R)).$
Thus we obtain $V(I)=\mathrm{DSupp}(S_{dg}(R))$, that is, the diagonal support scheme of $\mathrm{Hom}_{S_{dg}(R^{e})}(R,R)$ is $V(I)$. We thus completed the proof. ∎
@@
Second, since $R$ is a Gorenstein Noetherian algebra, $R^{e}$ is also a Gorenstein Noetherian algebra. Hence, $R^{e}$ admits a bounded injective $R^{e}$-module resolution, denoted by $J^{*}$. There is a quasi-isomorphism of $R^{e}$-modules complexes:
$\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),J^{*})\rightarrow\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),R^{e})$